A capacitor is charged to a potential of 12.0 V and is then connected to a voltmeter having an internal resistance of 3.10 M Ohms. After a time of 4.20 s the voltmeter reads 2.8 V.

A. What is the capacitance (F)?

B. What is the time constant of the circuit (s)?
Answer #1
Concepts and reason

The concept required to solve this problem is capacitance, the time constant, and the voltage drop across the capacitor.

Initially, find the voltage drop across the capacitor. Then, find the capacitance of the capacitor. Later, find the charge on the capacitor and the current through the capacitor. Finally, calculate the time constant.

Fundamentals

The expression for the voltage drop in the capacitor is as follows:

V=V0etRCV = {V_0}{e^{\frac{{ - t}}{{RC}}}}

Here, V0{V_0} is the potential to which the capacitor is charged, t is the time taken in reading the unit of voltmeter, R is the internal resistance, and C is the capacitance.

The expression for the time constant is as follows:

τ=RC\tau = RC

Here, R is the resistance and C is the capacitance.

(A)

The expression for the voltage drop in the capacitor is as follows:

V=V0etRCV = {V_0}{e^{\frac{{ - t}}{{RC}}}}

Substitute 2.8 V for V, 4.20 s for t, 12.0 V for V0{V_0} , and 3.10MΩ3.10{\rm{ M}}\Omega for C in the equation 2.8V=(12.0V)e4.20s(3.10MΩ)C2.8V12.0V=e4.20s(3.10(106Ω1MΩ))C\begin{array}{c}\\2.8{\rm{ V}} = \left( {12.0{\rm{ V}}} \right){e^{\frac{{ - 4.20{\rm{ s}}}}{{\left( {3.10{\rm{ M}}\Omega } \right)C}}}}\\\\\frac{{2.8{\rm{ V}}}}{{12.0{\rm{ V}}}} = {e^{\frac{{ - 4.20{\rm{ s}}}}{{\left( {3.10\left( {\frac{{{{10}^6}{\rm{ }}\Omega }}{{1{\rm{ M}}\Omega }}} \right)} \right)C}}}}\\\end{array}

Solve for C.

0.23=e4.20s(3.10(106Ω1MΩ))CC=9.31×107F\begin{array}{c}\\0.23 = {e^{\frac{{ - 4.20{\rm{ s}}}}{{\left( {3.10\left( {\frac{{{{10}^6}{\rm{ }}\Omega }}{{1{\rm{ M}}\Omega }}} \right)} \right)C}}}}\\\\C = 9.31 \times {10^{ - 7}}{\rm{ F}}\\\end{array}

(B)

Substitute 3.10MΩ3.10{\rm{ M}}\Omega for R and 9.31×107F9.31 \times {10^{ - 7}}{\rm{ F}} for C in the equation τ=RC\tau = RC .

τ=(3.10MΩ)(9.31×107F)=(3.10MΩ)(106Ω1MΩ)(9.31×107F)=2.89s\begin{array}{c}\\\tau = \left( {3.10{\rm{ M}}\Omega } \right)\left( {9.31 \times {{10}^{ - 7}}{\rm{ F}}} \right)\\\\ = \left( {3.10{\rm{ M}}\Omega } \right)\left( {\frac{{{{10}^6}{\rm{ }}\Omega }}{{1{\rm{ M}}\Omega }}} \right)\left( {9.31 \times {{10}^{ - 7}}{\rm{ F}}} \right)\\\\ = 2.89{\rm{ s}}\\\end{array}

Ans: Part A

The capacitance of the capacitor is equal to 9.31×107F9.31 \times {10^{ - 7}}{\rm{ F}} .

Part B

The time constant of the circuit is equal to 2.89 s.